OpenStudy (anonymous):

How much dinner dish, if the fish costs twice as much as the meat. He needs to order 164 meat selections and 24 fish selections. His budget is $4240.

OpenStudy (anonymous):

let m for meat f for fish f=2m 164m+24f=$4240 164m+24(2m)=$4240 164m+48m=$4240 212m =$4240 m= $4240/212 m=20 164(20)+24f=$4240 3280 +24f= $4240 f = (4240-3280)/24 f =40

OpenStudy (anonymous):

If u like my answer, click good answer,ty